Rosa works due north and her husband, Alan drives due east to work.They leave work at the same time. When Rosa is 3 miles from home the distance between them is 1 more mile than Alan's distance. How many miles are both?
Answer by jorel1380(3719)

Let n and n+1 be the distance Alan is from work and the distance he is from Rosa, respectively. Then:
3²+n²=(n+1)²
n²+9=n²+2n+1
2n=8
n=4
☺☺☺☺